Find the missing diagonal of the rhombus that has an area of 12 feet squared and on diagonal of 6ft
Answer by AFKeying(5)

The answer is 4 feet.
you find the answer by breaking the rhombus into four triangles.
the formula for the area of the triangle is A=1/2bh
Since there is four triangles now, the formula is A=4(1/2)bh
You only use half of the diagonal for each triangle (this is important to remember)
Plug in your values and simplify.
A=4(1/2)bh
12=4(1/2)3h
12=(2)3h
12=6h
2=h
but remember you only use half of the diagonal! so your answer you are looking for is actually 4 feet.
Don't forget to label your units.